OpenStudy (anonymous):

9b^3=7b-5(9b^3-2b)= I have no idea

OpenStudy (istim):

Ok, so first multiply (9b^3-2b) by 5.

OpenStudy (istim):

Now, subtract 9b^3 from both sides.

OpenStudy (anonymous):

9b^3=17b-45b^3

OpenStudy (anonymous):

so the first step would it be -45b^3=10b

OpenStudy (istim):

Why the equal sign? IT's still a part of its original equation.

OpenStudy (anonymous):

sorry it should be a + sign

OpenStudy (anonymous):

i got 9b^3-45b^3=3b which would simplify to -36b^3=3b

OpenStudy (istim):

Whoops sorry. I actually would't know; I can only do this on paper.

OpenStudy (istim):

What happened to 7b and 2b (which was in the brackets)?

OpenStudy (anonymous):

sorrrrrrrrrrrrrry no wounder i'm confused i wrote wrong problem it should be 9b^3+7b-5(9b^3-2b)

OpenStudy (istim):

Next time please just enter the formula correctly. The humour in the question is appreciated, but makes the question confusing.

OpenStudy (istim):

And is it 0=9b^3+7b-5(9b^3-2b)? That would make a lot more sense.

OpenStudy (istim):

Hello?

OpenStudy (anonymous):

Im sorry im flipping from page to page and wrote it wrong but how did you get that cuz i got -36b^3+17b

OpenStudy (istim):

No, no, yours sounds right. What I said it just the type of formula that it would be formatted at the beginning. I fully believe that your answer is probably correct. I myself can't check; sorry. I'm overwhelmed with work at my end.

OpenStudy (anonymous):

oh no problem i thought you ment it =0 i was like OMG what did i do wrong lol

OpenStudy (istim):

Ok. Well, good luck with your other questions then! I didn't really do much this round; nice work.